OpenStudy (anonymous):

Solve and graph the absolute value inequality: |4x + 1| ≤ 5 number line with closed dots on −1.5 and 1 with shading going in the opposite directions. number line with open dots on −1.5 and 1 with shading in between. number line with closed dots on −1 and 1 with shading in between. number line with closed dots on −1.5 and 1 with shading in between.

OpenStudy (anonymous):

| 4x + 1 | <= 5 -5 <= 4x + 1 <= 5 --- subtract 1 -5 - 1 <= 4x <= 5 - 1 -6 <= 4x <= 4 --- divide by 4 -6/4 <= x <= 4/4 -1.5 <= x <= 1 closed dots on -1.5 and 1, shading in between

OpenStudy (anonymous):

thank u !!

OpenStudy (anonymous):

anytime :)